OpenStudy (anonymous):

Rose Bought A New Hat When She Was On Vacation. The Following Formula Describes The Total Amount Rose Paid: t=c+cb Where T=Total Cost Rose Paid For The Hat C=Price Of The Hat B=Sales Tax (Percent) If The Hat Cost $13.21 And The Sales Tax Was 6%, How Much Did Rose Pay For The Hat ?

OpenStudy (anonymous):

@evergirl

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Can You Help ?

OpenStudy (anonymous):

srry i suck at math

OpenStudy (anonymous):

Okay Thanks Anyway

OpenStudy (anonymous):

@hartnn

Parth (parthkohli):

I can help.

Parth (parthkohli):

Wait... the formula ain't right

OpenStudy (anonymous):

Oh Sorry Wait Just A Minute ,

Parth (parthkohli):

Where's the "B"?

OpenStudy (anonymous):

Here You Go I Fixed It ,

Parth (parthkohli):

Oh, gotcha.

Parth (parthkohli):

You have the formula,\[\rm t = c + cb\]and you have \(\rm c = 13.21\) and \(\rm b = 6\). You have to find what \(\rm t\) is.

Parth (parthkohli):

Substituting \(\rm c = 13.21 \) and \(\rm b = 6\),\[\rm t = 13.21 + (13.21\times 6\%)\]

OpenStudy (anonymous):

I Got 14.0026

Parth (parthkohli):

\[13.21 +\left(13.21 \times {6 \over 100}\right)\]

Parth (parthkohli):

Looks good enough. Because we have to round to nearest tenths, we are left with $14.

OpenStudy (anonymous):

So She Paid $14 Dollars For The New Hat

Parth (parthkohli):

Yep.

OpenStudy (anonymous):

Can You Help Me With 3 More